OpenStudy (anonymous):

Find a vector that has the same direction as <−4, 6, 4> but has length 6.

OpenStudy (anonymous):

find the length of this vector then multiply by whatever you need to make is 6

OpenStudy (anonymous):

ok ill try that. wouldnt that be tedious becuase of the sqrt(68)

OpenStudy (anonymous):

i guess, but if you want it to be 6 multiply by \(\frac{6}{\sqrt{68}}\)

OpenStudy (anonymous):

haha thanks! :D

OpenStudy (anonymous):

or maybe \(\frac{3}{\sqrt{17}}\)

OpenStudy (anonymous):

yea same difference its trivial now. QED bro!

OpenStudy (anonymous):

qed